OpenStudy (anonymous):

what percent of 60 is 12?

jigglypuff314 (jigglypuff314):

12 / 60 = gives you a decimal multiply decimal you get by 100 then put % after it :P

OpenStudy (anonymous):

x percent of 60 = 12 x/100 * 60 = 12 x = 12 * 100/60 x = 20

OpenStudy (anonymous):

i still dont really understand

jigglypuff314 (jigglypuff314):

12 / 60 = 0.2 0.2 * 100 = 20 so 20%

OpenStudy (anonymous):

okay, now i get it. Thanks so much!

OpenStudy (jdoe0001):

\(\large {\begin{array}{llll} percent&quantity\\ l00& 12\\ x& 60 \end{array}\implies \cfrac{100}{x}=\cfrac{12}{60}}\)

OpenStudy (anonymous):

simplify the expression below. Show your work. -(-7y+12)

jigglypuff314 (jigglypuff314):

-(-7y+12) distributive property -1(-7y+12) (-1)(-7y) + (-1)(12) then multiply 7y - 12